OpenStudy (nikki_skellington):

The price of tomatoes at a farm is $0.54 per pound. Which equation can be used to determine c, the total price of n pounds of tomatoes? c = 0.54n c = 0.54 over n c = 0.54 + n c = 0.54 − n

OpenStudy (nikki_skellington):

is it a?

OpenStudy (anonymous):

c=54n

OpenStudy (nikki_skellington):

u sure

OpenStudy (anonymous):

yes

OpenStudy (nikki_skellington):

The graph below shows the prices of different numbers of boots at a store: A graph is shown. The values on the x axis are 0, 2, 4, 6, 8, and 10. The values on the y axis are 0, 25, 50, 75, 100, and 125. Points are shown on ordered pairs 0, 0 and 2, 25 and 4, 50 and 6, 75 and 8, 100. These points are connected by a line. The label on the x axis is Number of Boots. The title on the y axis is Price in dollars. Which equation can be used to determine p, the cost of b boots? p = 12.50b p = 25b b = 25p b = 12.50p

OpenStudy (phi):

a line in slope-intercept form is y= mx + b the x is along the x-axis (sideways) and the y is vertical axis (up/down) in your problem they say boots is on the sideways axis so should you expect p= m*b or b= m*p ? any idea ?

OpenStudy (nikki_skellington):

i think b= m*p

OpenStudy (nikki_skellington):

sorry i mean p= m*b right?

OpenStudy (phi):

yes p= m*b because b (for boots I assume) is "sideways" along the x-axis m is the slope. change in y divided by change in x or in this case change in p divided by change in b can you find the slope ?

OpenStudy (nikki_skellington):

ok thx :D